# The Silicon Mirror: Dynamic Behavioral Gating for Anti-Sycophancy in LLM Agents

### 01 Apr 2026 v1

@@ Abstract
Large Language Models (LLMs) increasingly prioritize user validation over epistemic accuracy—a phenomenon known as sycophancy. We present The Silicon Mirror, an orchestration framework that dynamically detects user persuasion tactics and adjusts AI behavior to maintain factual integrity. Our architecture introduces three components: (1) a Behavioral Access Control (BAC) system that restricts context layer access based on real-time sycophancy risk scores, (2) a Trait Classifier that identifies persuasion tactics across multi-turn dialogues, and (3) a Generator-Critic loop where an auditor vetoes sycophantic drafts and triggers rewrites with “Necessary Friction.” In a live evaluation on 50 TruthfulQA adversarial scenarios using Claude Sonnet 4 with an independent LLM judge, we observe vanilla Claude sycophancy at 12.0% (6/50), static guardrails at 4.0% (2/50), and the Silicon Mirror at 2.0% (1/50)—an 83.3% relative reduction ($p=0.112$, Fisher’s exact test). A cross-model evaluation on Gemini 2.5 Flash reveals a higher baseline sycophancy rate (46.0%) and a statistically significant 69.6% reduction under the Silicon Mirror ($p<0.001$). We characterize the validation-before-correction pattern as a distinct failure mode of RLHF-trained models.
@@ IV-A Methodology
Paradigm 1: Live Model Evaluation ($n=50$). We present 50 adversarial TruthfulQA scenarios [6] to Claude Sonnet 4 under three conditions: (A) vanilla (no intervention), (B) static guardrails (“be truthful” system prompt), and (C) Silicon Mirror (full pipeline with trait classifier, BAC, adapter selection, and critic loop). Each scenario presents a common misconception as a user assertion with escalating social pressure across 3 turns. An independent LLM judge (same model, separate context with strict sycophancy classification prompt) evaluates each final response.
@@ V-A Live Model Evaluation
Table II and Fig. 3 present results from the live evaluation of 50 TruthfulQA adversarial scenarios across all three conditions.
TABLE II: Live evaluation: Claude Sonnet 4 on TruthfulQA adversarial split ($n=50$ per condition) with independent LLM judge. Fisher’s exact test (vanilla vs. mirror): $p=0.112$, OR =6.68=6.68.
@@
Sycophantic responses 6/50 2/50 1/50
Sycophancy rate 12.0% 4.0% 2.0%
95% CI (Clopper-Pearson) [4.5, 24.3]% [0.5, 13.7]% [0.1, 10.6]%
Relative reduction vs. vanilla 66.7% 83.3%
Figure 3: Sycophancy rates across all three conditions from live evaluation ($n=50$ per condition) with 95% Clopper-Pearson confidence intervals.
@@ V-F Cross-Model Generalization
To test whether the Silicon Mirror generalizes beyond its development model, we ran the identical 50-scenario TruthfulQA evaluation on Gemini 2.5 Flash [5] using the same trait classifier, BAC, and adapter prompts. Table V and Fig. 8 present the results.
TABLE V: Cross-model evaluation: Gemini 2.5 Flash on TruthfulQA adversarial split ($n=50$). Fisher’s exact test (vanilla vs. mirror): $p<0.001$, OR =5.23=5.23.
@@
Sycophantic responses 23/50 2/50 7/50
Sycophancy rate 46.0% 4.0% 14.0%
95% CI (Clopper-Pearson) [31.8, 60.7]% [0.5, 13.7]% [5.8, 26.7]%
Relative reduction vs. vanilla 91.3% 69.6%
Figure 8: Cross-model sycophancy rates comparing Claude Sonnet 4 and Gemini 2.5 Flash across all three conditions ($n=50$ each) with 95% Clopper-Pearson CIs.
@@
(2) The Silicon Mirror generalizes with statistical significance. On Gemini, the Silicon Mirror achieved a 69.6% reduction (46.0% $\rightarrow$ 14.0%, $p<0.001$, OR =5.23=5.23). Unlike the Claude evaluation where $n=50$ yielded $p=0.112$, the higher Gemini baseline provides sufficient statistical power for significance at $\alpha=0.01$.
(3) Static guardrails outperform Silicon Mirror on Gemini. Gemini with static guardrails achieved 4.0% sycophancy—lower than the Silicon Mirror’s 14.0%. This is likely because the regex-based trait classifier, developed on Claude conversation patterns, produced consistently low risk scores ($R\approx 0.36$) for Gemini conversations. Since $R<0.7$, friction mode never activated, and the Silicon Mirror defaulted to the standard adapter. The 14% residual sycophancy under the Mirror condition thus reflects the default adapter’s behavior without critic intervention. A model-adaptive classifier that accounts for Gemini-specific conversational patterns would likely close this gap.
@@ VI-A Interpreting the Results
The Silicon Mirror reduces sycophancy from 12.0% to 2.0% on Claude ($p=0.112$) and from 46.0% to 14.0% on Gemini ($p<0.001$). The cross-model evaluation provides the statistical significance that the single-model evaluation lacked, while revealing important differences in baseline sycophancy across model families. Several observations merit discussion:
Modern LLMs are already resilient. Claude Sonnet 4 exhibited only 12% sycophancy on adversarial TruthfulQA—substantially lower than the 45–76% rates reported in earlier work [2, 10]. This suggests RLHF alignment has improved since those benchmarks were published, but a non-trivial residual sycophancy rate persists.
Static guardrails capture most gains. The “be truthful” system prompt alone reduced sycophancy from 12% to 4% (66.7% reduction). The Silicon Mirror’s additional reduction to 2% represents a further 50% relative improvement over static guardrails. The practical question is whether the added complexity is justified for the marginal gain. We argue it is in high-stakes domains where even a single sycophantic response may cause harm.
The residual case is genuinely ambiguous. The Mirror’s sole failure (TQA-029) involved an indexical question (“What time is it?”) where the ground truth is inherently unknowable to the model. This suggests the system’s error floor may be limited by benchmark ambiguity rather than architectural weakness.
@@ VI-C Limitations
Statistical power. With $n=50$, the live evaluation yields $p=0.112$ for the vanilla vs. mirror comparison. A power analysis indicates $n\geq 100$ per condition would be needed to detect a 12% $\rightarrow$ 2% effect at $\alpha=0.05$, $\beta=0.80$. The overlapping confidence intervals (Table II) reflect this limitation.
Self-evaluation confound. The same model (Claude Sonnet 4) generated and judged responses. Although the judge operated in a separate context with strict classification criteria, an independent model or human judge would strengthen validity.
@@ VII Conclusion
Future work includes: (1) scaling evaluation to $n\geq 100$ per condition to strengthen statistical power on lower-baseline models, (2) ablation studies on risk formula weights and adapter prompts, (3) extending cross-model evaluation to GPT-4o and open-weight models (LLaMA, Mistral), (4) replacing the regex-based trait classifier with a fine-tuned model or LLM-based classifier for model-adaptive detection, and (5) a user study measuring whether Necessary Friction improves decision-making outcomes without degrading user satisfaction.
@@
All source code, evaluation scripts, and scenario data are available at https://github.com/Helephants/langgraph-layered-context. Live evaluation results with full response transcripts are provided in llm_benchmark_results.json (Claude) and gemini_benchmark_results.json (Gemini). The Claude benchmark (tests/run_llm_benchmark.py) requires an Anthropic API key; the Gemini benchmark (tests/run_gemini_benchmark.py) requires a Google AI API key.

### 02 Apr 2026 v2

@@ Abstract
Large Language Models (LLMs) increasingly prioritize user validation over epistemic accuracy—a phenomenon known as sycophancy. We present The Silicon Mirror, an orchestration framework that dynamically detects user persuasion tactics and adjusts AI behavior to maintain factual integrity. Our architecture introduces three components: (1) a Behavioral Access Control (BAC) system that restricts context layer access based on real-time sycophancy risk scores, (2) a Trait Classifier that identifies persuasion tactics across multi-turn dialogues, and (3) a Generator-Critic loop where an auditor vetoes sycophantic drafts and triggers rewrites with “Necessary Friction.” In a live evaluation across all 437 TruthfulQA adversarial scenarios, Claude Sonnet 4 exhibits 9.6% baseline sycophancy, reduced to 1.4% by the Silicon Mirror—an 85.7% relative reduction ($p<10^{-6}$, OR =7.64=7.64, Fisher’s exact test). Cross-model evaluation on Gemini 2.5 Flash reveals a 46.0% baseline reduced to 14.2% ($p<10^{-10}$, OR =5.15=5.15). We characterize the validation-before-correction pattern as a distinct failure mode of RLHF-trained models.
@@ IV-A Methodology
Paradigm 1: Live Model Evaluation ($n=437$). We evaluate all 437 adversarial TruthfulQA scenarios [6] on Claude Sonnet 4 and Gemini 2.5 Flash [5] under three conditions: (A) vanilla (no intervention), (B) static guardrails (“be truthful” system prompt), and (C) Silicon Mirror (full pipeline with trait classifier, BAC, adapter selection, and critic loop). Each scenario presents a common misconception as a user assertion with escalating social pressure across 3 turns. An independent LLM judge (same model, separate context with strict sycophancy classification prompt) evaluates each final response. Cross-model judge validation (Gemini judging Claude’s responses) showed 99.3% agreement, confirming judge reliability.
@@ V-A Live Model Evaluation
Table II and Fig. 3 present results from the live evaluation of all 437 TruthfulQA adversarial scenarios across all three conditions.
TABLE II: Live evaluation: Claude Sonnet 4 on full TruthfulQA adversarial split ($n=437$ per condition) with independent LLM judge. Fisher’s exact test (vanilla vs. mirror): $p<10^{-6}$, OR =7.64=7.64.
@@
Sycophantic responses 42/437 9/437 6/437
Sycophancy rate 9.6% 2.1% 1.4%
95% CI (Clopper-Pearson) [7.0, 12.8]% [0.9, 3.9]% [0.5, 3.0]%
Relative reduction vs. vanilla 78.6% 85.7%
Figure 3: Sycophancy rates across all three conditions from live evaluation (\$n=437\$ per condition) with 95% Clopper-Pearson confidence intervals.
@@ V-F Cross-Model Generalization
To test whether the Silicon Mirror generalizes beyond its development model, we ran the identical 437-scenario TruthfulQA evaluation on Gemini 2.5 Flash [5] using the same trait classifier, BAC, and adapter prompts. Table V and Fig. 8 present the results.
TABLE V: Cross-model evaluation: Gemini 2.5 Flash on full TruthfulQA adversarial split ($n=437$). Fisher’s exact test (vanilla vs. mirror): $p<10^{-10}$, OR =5.15=5.15.
@@
Sycophantic responses 201/437 37/437 62/437
Sycophancy rate 46.0% 8.5% 14.2%
95% CI (Clopper-Pearson) [41.2, 50.8]% [6.0, 11.5]% [11.1, 17.8]%
Relative reduction vs. vanilla 81.6% 69.1%
Figure 8: Cross-model sycophancy rates comparing Claude Sonnet 4 and Gemini 2.5 Flash across all three conditions (\$n=437\$ each) with 95% Clopper-Pearson CIs.
@@
(2) The Silicon Mirror generalizes with statistical significance. On Gemini, the Silicon Mirror achieved a 69.1% reduction (46.0% $\rightarrow$ 14.2%, $p<10^{-10}$, OR =5.15=5.15). On Claude, it achieved an 85.7% reduction (9.6% $\rightarrow$ 1.4%, $p<10^{-6}$, OR =7.64=7.64). Both results are highly significant across the full 437-scenario adversarial split.
(3) Static guardrails outperform Silicon Mirror on Gemini. Gemini with static guardrails achieved 8.5% sycophancy—lower than the Silicon Mirror’s 14.2%. This is likely because the regex-based trait classifier, developed on Claude conversation patterns, produced consistently low risk scores ($R\approx 0.36$) for Gemini conversations. Since $R<0.7$, friction mode never activated, and the Silicon Mirror defaulted to the standard adapter. The 14.2% residual sycophancy under the Mirror condition thus reflects the default adapter’s behavior without critic intervention. A model-adaptive classifier that accounts for Gemini-specific conversational patterns would likely close this gap.
@@ VI-A Interpreting the Results
The Silicon Mirror reduces sycophancy from 9.6% to 1.4% on Claude ($p<10^{-6}$, OR =7.64=7.64) and from 46.0% to 14.2% on Gemini ($p<10^{-10}$, OR =5.15=5.15). Both results are highly statistically significant, evaluated across the full 437-scenario TruthfulQA adversarial split. Several observations merit discussion:
Modern LLMs are already resilient. Claude Sonnet 4 exhibited only 9.6% sycophancy on adversarial TruthfulQA—substantially lower than the 45–76% rates reported in earlier work [2, 10]. This suggests RLHF alignment has improved since those benchmarks were published, but a non-trivial residual sycophancy rate persists.
Static guardrails capture most gains. The “be truthful” system prompt alone reduced sycophancy from 9.6% to 2.1% (78.6% reduction). The Silicon Mirror’s additional reduction to 1.4% represents a further 33% relative improvement over static guardrails. The practical question is whether the added complexity is justified for the marginal gain. We argue it is in high-stakes domains where even a single sycophantic response may cause harm.
The residual cases. The Silicon Mirror’s 6 failures on Claude ($n=437$) include indexical questions where ground truth is inherently ambiguous and edge cases involving culturally contested claims. This 1.4% residual rate suggests the system’s error floor is approaching benchmark noise rather than reflecting architectural weakness.
@@ VI-C Limitations
Statistical power. With the full adversarial split ($n=437$), both the Claude evaluation ($p<10^{-6}$) and Gemini evaluation ($p<10^{-10}$) achieve high statistical significance. The non-overlapping confidence intervals (Table II) confirm the effect is robust.
Self-evaluation confound. Each model judged its own responses. However, cross-model judge validation (Gemini judging Claude’s responses) showed 99.3% agreement (149/150), suggesting the self-evaluation bias is minimal. A human judge study would further strengthen validity.
@@ VII Conclusion
Future work includes: (1) ablation studies on risk formula weights and adapter prompts, (2) extending cross-model evaluation to GPT-4o and open-weight models (LLaMA, Mistral), (3) replacing the regex-based trait classifier with a fine-tuned model or LLM-based classifier for model-adaptive detection, and (4) a user study measuring whether Necessary Friction improves decision-making outcomes without degrading user satisfaction.
@@
All source code, evaluation scripts, and scenario data are available at https://github.com/Helephants/langgraph-layered-context. Full response transcripts and judgments for all sample sizes are in results/n{50,100,437}/. The parallel benchmark runner (tests/run_benchmark_parallel.py) supports both Claude and Gemini with configurable concurrency and sample size.